Mr. Stanley mowed 3/4 of his lawn in 1 1/4 hours. Mr. Samuels, who works twice as fast, finished mowing the lawn. How many minutes did mr. Samuels work?
A)12 1/2
B)15
C)25
D)38
E)50

Answer by lwsshak3(11628)

Mr. Stanley mowed 3/4 of his lawn in 1 1/4 hours. Mr. Samuels, who works twice as fast, finished mowing the lawn. How many minutes did mr. Samuels work?
**
Mr Stanley's work rate=(3/4)/75=.75/75=.01=1/100 (which means he could have finished mowing the lawn by himself working 25 more minutes at this work rate)
..
Mr Samuel's work rate=twice that of Mr Stanley's=2*(1/100)=1/50
let x=minutes Mr Samuel worked to finish mowing the lawn
3/4+x*1/50)=100% completion of mowing the lawn
.75+.02x=1
.02x=1-.75=.25
x=.25/.02=12.5
ans:
Minutes Mr Samuel worked to finish mowing the lawn=12.5
A is the correct answer